Nov 7 10 1 6) Calculate the number of moles and the number of grams of the indicated solutes present in each if the following solution samples. a) 127 mL of 0.105 M HNO b) 155 ml of 15.1 M NH,
Answer #1

a) Number of moles of HNO3 = molarity * volume of solution in L

Number of moles of HNO3 = 0.105* 0.127 L = 0.0133 mole

Therefore, the number of moles of HNO3 = 0.0133 mole

numebr of moles of HNO3 = mass of HNO3 / molar mass of HNO3

0.0133 mole = mass of HNO3 / 63.01 g/mol

mass of HNO3 = 0.838 g

Therefore, the mass of HNO3 = 0.838 g

b) Number of moles of NH3 = molarity * volume of solution in L

Number of moles of NH3 = 15.1 * 0.155 L = 2.34 mole

Therefore, the number of moles of NH3 = 2.34 mole

number of moles of NH3 = mass of NH3 / molar mass of NH3

2.34 mole = mass of NH3 / 17.031 g/mol

mass of NH3 = 39.9 g

Therefore, the mass of NH3 = 39.9 g